Which of the following is not 1/3 of an integer?
F. -7
G. -1/3
H. 1
J. 1/3
K. 1/2
Please give a detailed and simple explanation!! Thanks

Answer by stanbon(75887)

Which of the following is not 1/3 of an integer?
F. -7
G. -1/3
H. 1
J. 1/3
K. 1/2
------
For each of A,B,C..., check the answer to (1/3)x = value.
Example::
(1/3)x = -7
x = -21, which IS an integer
-----
Check::
(1/3)x = -1/3
x = -1, which IS an integer
------
Check::
(1/3)x = 1/2
x = 3/2, which is NOT an integer
-----------------
Ans: 1/2
==============
Cheers,
Stan H.
-------------